OpenStudy (anonymous):

how do you solve -13=2b-b-10

OpenStudy (midhun.madhu1987):

-13+10=2b-b

OpenStudy (anonymous):

but after what do you do

Nnesha (nnesha):

just combine like terms -13 +10 = ?? 2b -b = ??

OpenStudy (midhun.madhu1987):

2b-b=?